When a certain gas under a pressure of 5.00 x 10^6 Pa at 25.08 oC is allowed to expand to 3.00 times its original volume, its final pressure is 1.07 x 10^6 Pa. What is its final temperature?
1
Expert's answer
2021-02-10T13:14:00-0500

Answer

Using idea gas equation

PV=nRT

So final temperature

"T_2=\frac{T_1P_2V_2}{P_1V_1}"

"=\frac{298.23\times 1.07\times10^6\times3V_1}{5\times10^6\times V_1}"

=191.36K